It takes a groundskeeper 40 minutes to prepare a softball field for a game. It takes his assistant 50 minutes to prepare the same field. How long will it take if they work together to prepare the field?
Answer by richwmiller(17219)

1/40+1/50=1/
x = 200/9
x = 22.222